TANGENTS AND NORMALS
find the points on the curve  y = x^3 – 2x^2  – 2x  at which the tangent lines are parallel to the line y=2x -3

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Hello Student,
Please find answer to your question below

Slope of line = 2
y = x^3 - 2x^2 - 2x
y' = 3x^2 - 4x - 2
3x^2 - 4x - 2 = 2
3x^2 - 4x - 4 = 0
(3x+2)(x-2) = 0
x= 2, \frac{-2}{3}
Put in the curve and get y.